If X=2+root3 the find x-1/x and x square-1/x square

If X=2+root3  the find x-1/x and x square-1/x
square

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
x = 2 + sqrt (3)
Now
1/x = 1 /(2 + sqrt3)
On rationalization
1/x = 2 - sqrt(3)
Hence
x - 1/x =  2 * sqrt(3)
Similarly
x + 1/x = 4
Now
x² - 1/x² = (x + 1/x)(x-1/x) = 8 * sqrt(3)